OpenStudy (kewlgeek555):

Lines p and q are parallel. Explain how to determine the measure of angle E. Two parallel lines, p and q, intersected by transversal t. The angles where q and t intersect are labeled A, B, C, and eighty-three degrees. The corresponding angles where p and t intersect are labeled D, E, F, and G.

OpenStudy (kewlgeek555):

I need help, Please? I don't know how to make a model or even answer the question!

OpenStudy (anonymous):

Wasn't there any kind of image or diagram included with that question?

OpenStudy (kewlgeek555):

@qweqwe123123123123111 no, there was not.

OpenStudy (anonymous):

Then you can use this one. I made it from your description. The angles ABCDEFG were rather ambiguously described, so I just laid them out in order.

OpenStudy (kewlgeek555):

Thank you! Based on the drawing you made, the answer would be that Angle E is equal to 83 degrees because is the same angle as the unknown angle?

OpenStudy (anonymous):

No. Adjacent angles along the same side of a line total 180 degrees. So A+C=180, A+B=180, etc. Angles diagonally opposite are equal, so that: E=F, D=G, etc Now here, E is congruent to B. Using the info I just gave you, what is the angle of B?

OpenStudy (anonymous):

Are you there?

OpenStudy (kewlgeek555):

Sorry.

OpenStudy (kewlgeek555):

OH! SO the E is equal to B. 180 - 83 = 97 97 is the measure of E. Thanks!